OpenStudy (anonymous):

Find a quadratic function to model the values in the table. Predict the value of y for x = 6. x | y -1 2 0 -2 3 10

jimthompson5910 (jim_thompson5910):

are you familiar with polynomial interpolation?

OpenStudy (anonymous):

no i am not.

jimthompson5910 (jim_thompson5910):

y = ax^2 + bx + c is your basic general quadratic equation When x = -1 and y = 2, we plug them in to get y = ax^2 + bx + c 2 = a(-2)^2 + b(-2) + c 2 = 4a - 2b + c 4a - 2b + c = 2 ------------------------------------------------------- When x = 0 and y = -2, we plug them in to get y = ax^2 + bx + c -2 = a(0)^2 + b(0) + c -2 = 0a + 0b + c -2 = 0 + 0 + c -2 = c c = -2 This means that 4a - 2b + c = 2 becomes 4a - 2b + (-2) = 2 which turns into 4a - 2b = 4

jimthompson5910 (jim_thompson5910):

When x = 3 and y = 10, we plug them in to get y = ax^2 + bx + c 10 = a(3)^2 + b(3) + c 10 = 9a + 3b + c 10 = 9a + 3b + (-2) 9a + 3b = 12

jimthompson5910 (jim_thompson5910):

So we now have this system of equations 4a - 2b = 4 9a + 3b = 12

jimthompson5910 (jim_thompson5910):

oh wait i made a typo I'm just catching now...

jimthompson5910 (jim_thompson5910):

When x = -1 and y = 2, we plug them in to get y = ax^2 + bx + c 2 = a(-1)^2 + b(-1) + c 2 = a - b + c a - b + c = 2 a - b + (-2) = 2 a - b = 2 + 2 a - b = 4

jimthompson5910 (jim_thompson5910):

So the system is a-b = 4 9a + 3b = 12 Solve this to find 'a' and 'b'

OpenStudy (anonymous):

a = 2 and b = -2?

jimthompson5910 (jim_thompson5910):

good, so a = 2, b = -2 and c = -2 So the quadratic equation is y = 2x^2 - 2x - 2

jimthompson5910 (jim_thompson5910):

Your last step is to plug in x = 6 and evaluate to find y.

OpenStudy (anonymous):

y=70-2x

jimthompson5910 (jim_thompson5910):

y = 2x^2 - 2x - 2 y = 2(6)^2 - 2(6) - 2 y = 2(36) - 2(6) - 2 y = 72 - 12 - 2 y = 60 - 2 y = 58

jimthompson5910 (jim_thompson5910):

So when x = 6, y is y = 58

OpenStudy (anonymous):

oh okay I got it. Thanks for helping!!!! :)

jimthompson5910 (jim_thompson5910):

np